OpenStudy (anonymous):

Find the probability of drawing an ace from a fair deck of 52 cards with a single draw. 4/13 1/13 1/4 1/52

OpenStudy (anonymous):

there are 4 aces and 52 total cards 4/52 = 1/13

OpenStudy (anonymous):

Find the probability of rolling the number five with a single toss of a fair six-sided die. 1/2 1/6 1/4 1/5 IS IT 1/6

OpenStudy (anonymous):

there are six different numbers you can get: 1, 2, 3, 4, 5, and 6. the probability of getting a 5 is 1/6. you're right.

OpenStudy (anonymous):

Find the probability of rolling a sum of seven with a single toss of two fair six-sided dice. 0 1/7 1/6 1/9

OpenStudy (anonymous):

IM THINKIN 1/6

OpenStudy (anonymous):

okay, 36 combinations total (6*6 = 36) you can roll 1 and 6 6 and 1 5 and 2 2 and 5 3 and 4 4 and 3 that's 6 possibilities. 6/36 = 1/6. you're right.

OpenStudy (anonymous):

Find the probability of getting all heads on a toss of three fair coins. 1/8 1/3 1/2 3/8

OpenStudy (anonymous):

1/2 is the prboability of getting heads once. 3 times, (1/2)(1/2)(1/2) = 1/8

OpenStudy (anonymous):

*probability

OpenStudy (anonymous):

If we toss three coins 240 times, how many times can we expect the coins to have three tails showing? 30 120 80 90

OpenStudy (anonymous):

I DONT KNOW THIS ONE

OpenStudy (anonymous):

(1/2)(1/2)(1/2) = 1/8 for 3 tails on 3 coins 1/8(240) = 30

OpenStudy (anonymous):

Simplify: 4! 24 6 120 240

OpenStudy (anonymous):

4! = 4(3)(2)(1) = 24